OpenStudy (wendyluan):

can someone help with this algebra problem? If F(x)=x(x)-1 and g(x)=1-2x find the value of: 1) g(a+1)-g(a) 2) g(1/f3)

OpenStudy (anonymous):

what this F(x) has to do here

OpenStudy (wendyluan):

g(a+1)= 1-2(a+1) =1-2a-2 =-2a-1 g(a)=1-2a g(a+1)-g(a)= -2a-1-(1-2a) =-2a-1-1+2a =-2 ============================================